OpenStudy (anonymous):

If f(x) = x + 3, find the value of x if f(x) = 9. 6 12 3 9

OpenStudy (anonymous):

f(x) = x + 3=9 solve for x what will u get?

OpenStudy (anonymous):

Thank youuuuu (:

OpenStudy (anonymous):

given is f(x)=9 problem wants to find a value for x for which x+3=9

OpenStudy (anonymous):

x+3=9 then x=6

OpenStudy (jiteshmeghwal9):

sorry @mukushla yes u r right:)

OpenStudy (anonymous):

no problem